OpenStudy (anonymous):

Determine whether each series is convergent or divergent. (1/5+1^2) + (1/5+2^2) + (1/5+3^2)

OpenStudy (anonymous):

is it pretty clear that \[2^2,3^2,4^2,...\] gets bigger and bigger?

OpenStudy (anonymous):

oh maybe i am reading it incorrectly

OpenStudy (anonymous):

i know an= 1/5+n^2

OpenStudy (anonymous):

i want to use the ratio test.

OpenStudy (anonymous):

\[a_n=\frac{1}{5+n^2}\]?

OpenStudy (anonymous):

yeah thats what i meant

OpenStudy (anonymous):

i wouldn't bother with the ratio test \[\frac{1}{5+n^2}<\frac{1}{n^2}\]

OpenStudy (anonymous):

the one on the right is well known to converge in fact, it converges to the unlikely number \(\frac{\pi}{6}\) but you don't need to use that, just that is ia a p series or whatever they are called

OpenStudy (anonymous):

its cuz i didnt know how to use the comparison test

OpenStudy (anonymous):

in simpler english, comparison test is enuf

OpenStudy (anonymous):

idk what that is

OpenStudy (anonymous):

compare the terms to a series that you know converges if they are less, then that series converges as well

OpenStudy (anonymous):

it is the simplest test to use requires no algebra no taking ratios no taking roots just comparing

OpenStudy (anonymous):

conversely if each term is larger than a series you know diverges, then your series diverges as well

OpenStudy (anonymous):

ok thnx, and to think i was going to use the long way lolz

OpenStudy (anonymous):

make your life as simple as possible, math is complicated as it is

OpenStudy (anonymous):

ikr xp

OpenStudy (anonymous):

but i still love it XD

OpenStudy (anonymous):

lol me too

OpenStudy (anonymous):

man if you are really in high school, you must really love it !!

OpenStudy (anonymous):

yeah im in high school, a junior

OpenStudy (anonymous):

you will have your PhD by the time you are 21

OpenStudy (anonymous):

that'll be awesome